# FinAudio: A Benchmark for Audio Large Language Models in Financial Applications

### v1

@@ 5.2 Results on summarization task
Dataset Whisper Qwen2-Audio-7B-Instruct Gemini-1.5-flash Rouge-L BertScore Rouge-L BertScore Rouge-L BertScore FinAudioSum 0.053 0.514 0.048 0.467 0.072 0.553
Table 3: Summarization performance using Rouge-L and BertScore metrics.
@@
Given the high WER in the ASR tasks, effective summarization cannot be reliably performed by models with excessive transcription errors. Therefore, we excluded the SALMONN model from the summarization evaluation. Balancing model performance and computational cost, we selected three models—Whisper-v3, Qwen2-Audio-7B-Instruct, and Gemini-1.5-flash for subsequent summarization experiments.
@@
5.3 Prompt Robustness Analysis
The previous experiments indicated that different models respond differently to the prompt. Therefore, we further conducted a prompt robustness analysis on three AudioLLMs: Whisper-v3, Qwen2-Audio-7B, and Qwen2-Audio-7B-Instruct - across three ASR datasets. Following the approach from AudioBench Wang et al. (2024), we constructed an instruction set comprising 10 prompts. These prompts convey similar instructions expressed in diverse formulations (see Appendix B). For each audio input, we randomly selected one prompt from this set to observe how variations in prompts affected model performance, as presented in Figure 5.
@@ 5.3 Prompt Robustness Analysis
We compare the WER in Figure 5 using two prompt types: the random prompt described above and the fixed prompt from Section 4, used in the main results (Table 2). Results for Qwen2-7B-Instruct and Whisper-v3 remain consistent, indicating their robustness across prompts. However, the performance of Qwen2-7B notably declines under varied prompts, exhibiting higher WER. The substantial performance gap between Qwen2-7B and Qwen2-7B-Instruct indicates that this AudioLLM is sensitive to prompt variations, providing empirical support for the necessity of instruction-tuning in AudioLLM training.
5.4 Error Analysis
During the ASR tasks, we noticed that AudioLLMs sometimes misinterpret numerical information and specialized financial terms in speech. Such errors significantly impact the factual accuracy of AudioLLMs within financial contexts. Specific examples of these errors are provided in the Appendix C.
6 Conclusion and Research Outlook
We propose FinAudio, an AudioLLM Benchmark for financial audio data. We constructed a total of 400h+ of audio data for three tasks and evaluated it on seven AudioLLMs. Future research directions for applying AudioLLMs in the financial industry include (1) extending the input context window length of AudioLLM, (2) enhancing the models’ comprehension of numerical transcription and specialized financial terminology, and (3) improving their reasoning capabilities in financial contexts.
Limitations
First, due to the limited availability of financial audio data, FinAudio primarily focuses on ASR-related tasks. Second, FinAudio does not address financial audio scenarios involving languages other than English.
Ethics Statement
All data sources used in FinAduio are public. The authors take full responsibility for the development of FinAduio, ensuring that the publicly available part of the dataset does not contain personal information and conforms to established ethical guidelines.
Appendix A Dataset Source
In section 3.2, we outline the statistical information for the datasets used in evaluation. Below, we provide a detailed description of the original datasets.
There are two datasets consisting of short financial audio clips:
  • MDRM Qin and Yang (2019): This is the first dataset to use multimodal financial data for stock volatility and movement prediction. MDRM includes both audio recordings and text transcripts from the 2017 earnings calls of 500 major companies listed on the S&P 500 and traded on U.S. stock exchanges. There are a total of 280 companies with 572 earnings conference calls. The author further segmented earnings conference calls into individual sentences based on the text and conducted corresponding audio segments. There are a total of 88,829 audio clips. We split the dataset into training and test sets using a 75%–25% ratio, with the test portion utilized as the evaluation set for the benchmark.
  • SPGISpecch O’Neill et al. (2021): The dataset comprises 5,000 hours of earnings conference call recordings. The original audio is segmented by sentence and professionally transcribed into a structured format. Each clip is between 5s and 15s long. The SPGISpeech dataset is originally split into training and test subsets; we utilize only the test subset for evaluation.
The combined test sets from the two datasets described above contain a total of 61,549 clips. This scale is sufficient for the evaluation of model performance. The corresponding training sets can be utilized for further supervised fine-tuning of AudioLLMs.
Additionally, there are two other publicly available earnings conference call datasets containing raw audio recordings, which can be leveraged for the long financial audio ASR task.
  • Earnings-21 Del Rio et al. (2021): The dataset comprises 44 complete earnings conference call recordings from 2021. A distinguishing characteristic of Earnings-21 is its emphasis on named entities, financial jargon, and numerical data, which often pose difficulties for ASR models. The aligned transcripts are further enhanced with detailed formatting and entity tags, making it possible to conduct granular error analysis.
  • Earnings-22 Del Rio et al. (2022): This is a updated version of Earnings-21. It has 125 audio samples with a total of 120 hours.
Both datasets were utilized for evaluating model performance.
Appendix B Instruction Set
As described in Section 5.3, we established an instruction set comprising 10 distinct prompts to evaluate the robustness of AudioLLMs. These prompts convey the same meaning. The detailed prompts are listed as follows:
  • (1) Convert the audio speech into a text transcript.
  • (2) Transcribe the spoken words into written form.
  • (3) Listen to the audio and provide the text version.
  • (4) Transform the speech into a text document.
  • (5) Capture the spoken language and convert it to text.
  • (6) Decode the audio and give me the written transcription.
  • (7) Recognize the verbal communication and transcribe it into text.
  • (8) Turn the speech input into a text transcription.
  • (9) Process the audio speech and provide the text output.
  • (10) Translate the spoken conversation into written text.
To assess the robustness of AudioLLMs to variations in input instructions, we randomly select a instruction from the above instruction set for each audio input. This approach closely simulates real-world scenarios, where user commands often differ in wording yet convey the same intent.
Appendix C Error Analysis
The experimental results demonstrate that ASR performance significantly influences downstream tasks, such as summarization. Through detailed analysis of the outputs, we identified the following major error categories:
  • Financial terminology error: Financial audio frequently contains specialized financial proper nouns, leading AudioLLMs to produce translation errors. For example, the company name “NextEra Energy" was incorrectly transcribed as “Era Energy."
  • Numerical information error : Financial audio typically includes extensive numerical information, such as monetary amounts. During transcription, AudioLLMs frequently encounter issues like digit inconsistencies or missing monetary units.

### v2

@@ Appendix D Results on summarization task
Dataset Whisper Qwen2-Audio-7B-Instruct Gemini-1.5-flash GPT-4o-audio Rouge-L BertScore Rouge-L BertScore Rouge-L BertScore Rouge-L BertScore FinAudioSum 0.053 0.514 0.048 0.467 0.072 0.553 0.063 0.508
Table 4: Summarization performance using Rouge-L and BertScore.
@@
Given the high WER in the ASR tasks, effective summarization cannot be reliably performed by models with excessive transcription errors. Therefore, we excluded the SALMONN model from the summarization evaluation. Balancing model performance and computational cost, we selected three models—Whisper-v3, Qwen2-Audio-7B-Instruct, and Gemini-1.5-flash for summarization experiments.
@@
Appendix E Prompt Robustness Analysis
The previous experiments indicated that different models respond differently to the prompt. Therefore, we further conducted a prompt robustness analysis on three AudioLLMs: Whisper-v3, Qwen2-Audio-7B, and Qwen2-Audio-7B-Instruct - across three ASR datasets. Following the approach from AudioBench [wang2024audiobench], we constructed an instruction set comprising 10 distinct prompts to evaluate the robustness of AudioLLMs. These prompts convey similar instructions expressed in diverse formulations. The detailed prompts are listed as follows.
Number Prompt Set 1 Convert the audio speech into a text transcript. 2 Transcribe the spoken words into written form. 3 Listen to the audio and provide the text version. 4 Transform the speech into a text document. 5 Capture the spoken language and convert it to text. 6 Decode the audio and give me the written transcription. 7 Recognize the verbal communication and transcribe it into text. 8 Turn the speech input into a text transcription. 9 Process the audio speech and provide the text output. 10 Translate the spoken conversation into written text.
Table 5: The evaluation instruction set for AudioLLMs
For each audio input, we randomly selected one prompt from this set to observe how variations in prompts affected model performance. This approach closely simulates real-world scenarios, where user commands often differ in wording yet convey the same intent.
@@ Appendix E Prompt Robustness Analysis
We compare the WER in Figure 4 using two prompt types: the random prompt described above and the fixed prompt from Section C, used in the main results (Table 2). Results for Qwen2-7B-Instruct and Whisper-v3 remain consistent, indicating their robustness across prompts. However, the performance of Qwen2-7B notably declines under varied prompts, exhibiting higher WER. The substantial performance gap between Qwen2-7B and Qwen2-7B-Instruct indicates that this AudioLLM is sensitive to prompt variations, providing empirical support for the necessity of instruction-tuning in AudioLLM training.
Appendix F Error Analysis
The experimental results demonstrate that ASR performance significantly influences downstream tasks, such as summarization. Therefore, we take an error analysis. During the ASR tasks, we noticed that AudioLLMs sometimes misinterpret numerical information and specialized financial terms in speech. Such errors significantly impact the factual accuracy of AudioLLMs within financial contexts. Through detailed analysis of the outputs, we identified the following major error categories: 1) Financial terminology error: Financial audio frequently contains specialized financial proper nouns, leading AudioLLMs to produce translation errors. For example, the company name “NextEra Energy" was incorrectly transcribed as “Era Energy." 2) Numerical information error: Financial audio typically includes extensive numerical information, such as monetary amounts. During transcription, AudioLLMs frequently encounter issues like digit inconsistencies or missing monetary units.
This analysis will guide future model training by prioritizing improvements in numerical recognition and enhancing the model’s ability to accurately transcribe financial terminology.